### v4

# Synergizing Large Language Models and Task-specific Models for Time Series Anomaly Detection

Abstract
In anomaly detection, methods based on large language models (LLMs) can incorporate expert knowledge by reading professional document, while task-specific small models excel at extracting normal data patterns and detecting value fluctuations from training data of target applications. Inspired by the human nervous system, where the brain stores expert knowledge and the peripheral nervous system and spinal cord handle specific tasks like withdrawal and knee-jerk reflexes, we propose CoLLaTe, a framework designed to facilitate collaboration between LLMs and task-specific models, leveraging the strengths of both models for anomaly detection. In particular, we first formulate the collaboration process and identify two key challenges in the collaboration: (1) the misalignment between the expression domains of the LLMs and task-specific small models, and (2) error accumulation arising from the predictions of both models. To address these challenges, we then introduce two key components in CoLLaTe: a model alignment module and a collaborative loss function. Through theoretical analysis and experimental validation, we demonstrate that these components effectively mitigate the identified challenges and achieve better performance than both LLM-based and task-specific models.
1 Introduction
Recently, methods based on Large Language Models (LLMs) have demonstrated the capablility of reading professional documents, effectively acquiring human knowledge to guide relevant tasks. However, they are often insensitive to the value fluctuations in time series data, and their natural language processing-based representations do not align well with the characteristics of time series data Jin et al., (2024). In contrast, task-specific small scaled models, such as task-specific anomaly detection models (TSADM), typically lack the capabilities of reading professional documents and extracting expert knowledge. However, these models are specifically designed to learn patterns for anomaly detection and often exhibit superior performance when applied to well-matched anomaly detection datasets Zhou et al., (2023). Despite their strengths, TSADMs also have notable limitations. First, for complex systems, such as cloud service system and flight systems, which require lots of expert knowledge to detect and diagnose anomalies accurately, the anomaly assumptions of unsupervised anomaly detection methods can be inconsistent with these scenarios Zha et al., (2020). For instance, anomalies identified based on the voting mechanism of redundant aircraft channels Rice and Mccorkle, (1979) may differ from the outlier assumptions utilized by unsupervised anomaly detection methods Blázquez-García et al., (2021). Thus, researchers need to modify TSADMs to incorporate the background expertise to achieve optimal performance. For instance, anomaly detection methods tailored for cloud service monitoring Ma et al., (2021); Chen et al., 2024b or aircraft monitoring e Silva and Murcca, (2023) have been modified to suit these specific contexts. Second, in many practical scenarios, such as aircraft monitoring Nanduri and Sherry, (2016), it is not feasible to collect system monitoring data for all possible flight conditions. These conditions often represent distinct distributions and normal patterns in the monitoring data. This insufficient coverage of comprehensive data poses a significant challenge, greatly impeding the performance of TSADM.
Inspired by the human nervous system, which relies on the brain to store expert knowledge and extract general principles, while using the peripheral nervous system and spinal cord for specific tasks like the withdrawal reflex and knee-jerk reflex, we propose a framework called CoLLaTe. This framework aims to facilitate effective Collaboration between a Large Language model (analogous to the brain) and a Task-specific model (analogous to the peripheral nervous system and spinal cord) to leverage their complementary strengths for the anomaly detection task. By enabling a strong synergy between an LLM and a TSADM, we can conveniently embed expert domain knowledge by adjusting the prompt of LLM, including necessary background knowledge and eliminating the need to design specialized models for different application scenarios. This approach also mitigates performance degradation caused by insufficient monitoring data across diverse operational conditions, because LLMs excel at utilizing professional documents to incorporate domain knowledge, which is often expressed in natural language. Such knowledge is highly condensed and can cover more general situations compared to concrete data samples. For instance, compared with providing different time series and describing whether they are trigonometric or not, giving the rule $a\sin(w_{1}x+t_{1})+b\cos(w_{2}x+t_{2})$ is more condensed and can represent more general situations. Thus, in the cases where training data samples are scarce, LLMs can extract general standards or patterns from professional documents to guide anomaly detection and mitigate the performance degradation on unseen distributions.
To this end, CoLLaTe integrates an LLM, enhanced by professional documents, with a TSADM to assess the severity of anomalies for each time slot. Subsequently, CoLLaTe employs a conditional network to synthesize the judgments from the LLM and the TSADM, using the data representation from the TSADM as a condition, to produce a unified anomaly score. As illustrated in Fig. 1(a), the collated anomaly score retains the true positive judgments from both the LLM and the TSADM while reconciling false positives. During this collaborative process, two main challenges arise:
  • Misalignment between the expression domains of the LLM and the TSADM. The LLM and the TSADM interpret anomaly scores differently, meaning they may use the same score to represent different levels of anomaly severity. In the example shown in Fig. 1(b), most anomaly scores produced by the LLM are below 0.3, whereas the anomaly scores generated by the TSADM are predominantly centered around 0.4 after normalization. Consequently, an anomaly score of 0.4 might indicate a moderate anomaly for the LLM, but for the TSADM, it could signify a normal condition. Such inconsistencies in score interpretation can disrupt effective collaboration between the two models.
  • Prediction error accumulation. Both the LLM and the TSADM are subject to prediction errors. During the collaboration process, these errors may not cancel out but instead accumulate. Through theoretical analysis and experimental validation, we show that when using classical loss functions, such as Mean Squared Error (MSE), the errors from the two models tend to either compound or settle at a compromise between the higher and lower values, rather than canceling each other out.
To enable effective synergy between the LLM and the TSADM while addressing the challenges mentioned above, we make the following key contributions:
  • The CoLLaTe framework. We first formalize the collaboration process between the LLM and the TSADM, identifying the key challenges that arise during this process. To address these challenges, we propose a framework that facilitates seamless collaboration between the LLM and the TSADM.
  • LLM and TSADM alignment. We introduce a novel alignment module to harmonize the differing interpretations of anomaly scores between the LLM and the TSADM.
  • Theoretically sound collaborative loss function. We conduct a theoretical analysis of the aforementioned error accumulation phenomenon and propose a novel provably collaborative loss function to mitigate it.
  • We conduct extensive experiments to validate the effectiveness of CoLLaTe and each of its proposed modules.
(a) The performance of CoLLaTe, LLM and TSADM
(b) Misalignment between anomaly scores of the LLM and TSADM
(c) Anomaly score distribution of large language model
(d) Misalignment between original anomaly scores of LLM and TSADM
Figure 1: (a) The anomaly scores of LLM, TSADM, and CoLLaTe. (b) The normalized anomaly score distribution of TSADM and the fitted curve of anomaly scores from the LLM on Mustang Amvrosiadis et al., (2018). (c) shows the histogram of the anomaly score of LLM on the Mustang Amvrosiadis et al., (2018) dataset and the fitted curve of the anomaly score. (d) The original anomaly score distribution of TSADM for anomaly detection and the fitted curve of anomaly scores from the LLM on Mustang.
2 Method
2.1 Overview
The architecture of CoLLaTe is depicted in Fig. 2. The anomaly scores, $\dot{s}$ and $S$, are generated by a TSADM and an LLM, respectively, where a set-up pitch prompt is utilized to embed expert knowledge from professional document to the LLM (Appendix. A.12). The TSADM is an anomaly attention Xu et al., (2022) based anomaly detection model, whose architecture is illustrated in Appendix. A.10 in detail. These anomaly scores, $\dot{s}$ and $S$, typically represent different score interpretations, as the LLM and the TSADM can assign the same score to denote varying degrees of anomaly. To address this, we introduce an alignment module to align the different interpretations of the anomaly scores between the LLM and the TSADM. Following alignment, we employ a conditional network to synthesize the aligned scores from the TSADM and anomaly scores from the LLM, using the data representation $R$ as a condition. This process yields the collated anomaly score, $\hat{S}$, which is deemed as the final anomaly score to distinguish anomaly. Additionally, we propose a collaborative loss function that effectively leverages the complementary strengths of the LLM and the TSADM. Notably, we prove this loss function can mitigate error accumulation during the collaborative process between the LLM and the TSADM mathematically and experimentally.
Figure 2: The model architecture of CoLLaTe
2.2 Alignment Module
TSADM and large language models (LLMs) often interpret anomaly scores differently. In other words, they may assign the same anomaly score to represent varying degrees of anomaly severity. This discrepancy arises due to two main factors: value range misalignment and distribution misalignment. The value range misalignment is shown in Fig. 1(d), where we shift the fitted curve of LLM anomaly scores to the histogram of TSADM anomaly scores. Most reconstruction-based TSADMs use reconstruction error as the anomaly score, resulting in a highly variable range of scores. In contrast, the anomaly scores produced by LLMs are typically constrained to the range $[0,1]$. This problem can be easily solved by min-max normalization and we obtain the scaled anomaly scores of TSADM, $s$. This process is illustrated in detail in Appendix. A.13.
Even after scaling the anomaly scores of the TSADM, differences in the distribution of anomaly scores can still lead to inconsistent interpretations. As shown in Fig. 1(c), we use a half-Gaussian distribution, which will be elaborated in the following, to fit the histogram of LLM anomaly scores. Then, we shift the fitted curve to the histogram of the anomaly scores from the TSADM in Fig. 1(b), where we observe that most LLM anomaly scores fall below 0.3, whereas the TSADM’s anomaly scores are concentrated in the range of [0.3, 0.4]. As a result, an anomaly score of 0.4 might indicate moderate anomaly risks for the LLM, while the TSADM may interpret it as normal. In such cases, direct collaboration between the TSADM and the LLM without proper alignment could misrepresent their intended outputs. For example, if the LLM assigns a score of 0.4 to a time slot, the TSADM might incorrectly assume the LLM considers the time slot normal. To address this challenge, we propose an Alignment Module to align their interpretations and ensure effective collaboration.
Inspired by university grading policies, which align scores across professors by setting proportions for each grade level (e.g., controlling pass and excellence rates), we design the Alignment Module to align the distributions of anomaly scores from the LLM and the TSADM. This ensures that the same anomaly score represents a similar degree of anomaly for both models. To achieve this, we first fit the LLM’s anomaly score distribution $f(S)$ using a half-Gaussian distribution, as described in Eq.1. The fitting process involves creating a mirrored anomaly score set by reflecting the initial anomaly scores across the negative axis. This mirrored anomaly score set is then merged with the original anomaly score set, and the standard deviation of the combined set is estimated to determine $\sigma$ in Eq. 1.
$f(S)=\left\{\begin{matrix}\frac{2}{\sigma\sqrt{2\pi}}e^{-\frac{S^{2}}{2\sigma}},&S\geq 0\\ 0,&S<0\end{matrix}\right.$ (1)
Subsequently, we use a mapping function $\mathcal{M}(s)$ to map $s\in\{s_{i}\}_{i=0}^{n}$ from the TSADM to a new set, whose distribution should approach $f(S)$. To achieve this, we divide the value range of anomaly scores, $[0,1]$, into N bins. We use $c_{i}$ to denote how many mapped scores $\mathcal{M}(s)$ fall into the $i^{th}$ bin. Then, we minimize cross entropy to decrease the distance between $f(S)$ and the distribution mapped set as Eq. 2 shows, where $\hat{\mu},\hat{\sigma}$ are the mean and standard variance of distribution $f(S)$, $\mu_{M}=\frac{1}{n}\sum_{k=1}^{n}\mathcal{M}(s_{k})$, and $\hat{\lambda}_{1}$ and $\hat{\lambda}_{2}$ are hyperparameters and bigger than zero. One problem with Eq. 2 is that the counting process of $c_{i}$ is non-differentiable and could not be used in backforward propagation.
[…]
To solve the above problem, we prove that when $N$ approaches infinite, the objective function in Eq. 2 is equivalent to $-\frac{1}{n}\sum_{i=1}^{n}\log(f(\mathcal{M}(s_{i})))+\hat{\lambda}_{1}(\frac{1}{n}\sum_{i=1}^{n}\mathcal{M}(s_{i})-\hat{\mu})^{2}$, which is differentiable. The proof is given in Appendix. A.1. Consequently, we use Eq. 3 as our loss function.
$\min_{\mathcal{M}}.\mathcal{L}_{a}=-\frac{1}{n}\sum_{i=1}^{n}\log(f(\mathcal{M}(s_{i})))+\hat{\lambda}_{1}(\frac{1}{n}\sum_{i=1}^{n}\mathcal{M}(s_{i})-\hat{\mu})^{2}\\ +\hat{\lambda}_{2}(\frac{1}{n-1}\sum_{i=1}^{n}(\mathcal{M}(s_{i})-\mu_{M})^{2}-\hat{\sigma}^{2})^{2}$ (3)
2.3 Collaborative Loss
As mentioned before, we introduce a conditional network to synthesize the assessment of the TSADM and LLM. This network takes the aligned anomaly scores from the two models as input, along with the input data representation $R$ obtained from the TSADM as a condition. It outputs a collated anomaly score, as shown in Eq. 4, where $\mathcal{P}$ represents the trainable parameters of the conditional network. To optimize the collated anomaly scores in an unsupervised training process, we propose a collaborative loss function to leverage the complementary strength of LLM and TSADM.
Table 1: The average F1 score of TSADM and GPT4 on a aircraft monitoring dataset.
Contextual Point
TSADM 0.561 0.061
GPT4 0.409 0.231
$\hat{S}=\operatorname{ConditionalNet}(\mathcal{M}(s),S,R;\mathcal{P})$ (4)
Many anomaly detection studies categorize anomalies into contextual anomalies and point anomalies Schmidl et al., (2022); Tuli et al., (2022). We conducted experiments on two aircraft monitoring datasets, which are detailed in Appendix A.7. The results, presented in Tab. 1, demonstrate the complementary performance of the TSADM and the LLM in detecting contextual and point anomalies. Specifically, the F1 score for contextual anomalies is higher for the TSADM compared to GPT-4, while GPT-4 achieves a better F1 score for point anomalies than the TSADM. According to this insight, we propose a collaborative loss function that adaptively optimizes the collated anomaly score. Specifically, we begin by dividing the time series into patches Nie et al., (2023). For each time slot, we compute the average distance between the present time slot and all other time slots within the same patch, denoted as $D_{intra}$. Next, we calculate the average distance between the patch of the current time slot and all other patches, denoted as $D_{inter}$. Since anomalous time slots differ significantly from normal ones, while normal time slots tend to share similarities, the metrics $D_{intra}$ and $D_{inter}$ behave differently for point and contextual anomalies. For point anomalies, $D_{intra}$ should be large because the point anomaly stands out distinctly from the surrounding normal time slots. Conversely, $D_{inter}$ for point anomalies should be smaller, as the normal time slots within the current patch resemble those in other patches, reducing the average difference. For contextual anomalies, which typically involve a sustained deviation, the anomalies within an anomalous event are similar to each other but differ significantly from normal time slots. As a result, $D_{intra}$ should be small, while $D_{inter}$ should be large. Considering LLM can handle point anomaly better and TSADM can handle contextual anomaly better, we design the collaborative loss function in Eq. 5, where $\mathcal{L}(a,b)$ represents a loss function that measures the distance between $a$ and $b$.
$\mathcal{L}_{\beta}=\frac{D_{intra}}{D_{intra}+D_{inter}}\mathcal{L}(s,\hat{S})+\frac{D_{inter}}{D_{intra}+D_{inter}}\mathcal{L}(S,\hat{S})$ (5)
If we choose some popular loss functions as $\mathcal{L}(a,b)$, such as $\mathcal{L}(a,b)=(a-b)^{2}$, $\mathcal{L}_{\beta}$ will accumulate the prediction error of TSADM and LLM in the process of gradient descent, as shown in Theorem 1, which is proven in Appendix. A.2.
Assumption 1. The anomaly score prediction error of the TSADM is $\epsilon_{s}$, where $\epsilon_{s}$ obeys an unknown distribution $\mathcal{D}_{s}(\mu_{s},\sigma_{s})$, $\mu_{s}\neq 0$.
Assumption 2. The anomaly score prediction error of the LLM is $\epsilon_{S}$, which obeys an unknown distribution $\mathcal{D}_{S}(\mu_{S},\sigma_{S})$, $\mu_{S}\neq 0$.
Assumption 3. $\mathcal{L}_{\beta}^{*}$ in Eq. 6 is $L$-Lipschitz continuous.
$\begin{split}\mathcal{L}_{\beta}^{*}=-\sum_{i=1}^{n}\sum_{j=1}^{n}[(y_{i}-y_{j})(\hat{S}_{i}-\hat{S}_{j})]\end{split}$ (6)
We examined the validity of these Assumptions in the context of our experiment through empirical discussion, which is shown in Appendix. A.17.
Theorem 1. Given Assumption 1 - Assumption 2, when $\mathcal{L}(a,b)=(a-b)^{2}$, the difference between the optimal solution $\hat{S}^{*}$ of loss function $\mathcal{L}_{\beta}$ and ground truth $y$ is $\mathbb{E}[(\hat{S}^{*}-y)^{2}]\geq(\lambda_{1}\mu_{s}+\lambda_{2}\mu_{S})^{2}$, where $\lambda_{1}=\frac{D_{intra}}{D_{intra}+D_{inter}}$ and $\lambda_{2}=\frac{D_{inter}}{D_{intra}+D_{inter}}$.
From Theorem 1, we observe that when MSE is used as $\mathcal{L}$, the expected prediction error of the conditional network exceeds the weighted sum of the expected prediction errors of the TSADM and LLM. Given that the weights $\lambda_{1}$ and $\lambda_{2}$ are both greater than 0 and sum to 1, it indicates that the collaborative error, when using MSE as $L_{(a,b)}$, consistently exceeds the smaller of the errors produced by the LLM and TSADM. This phenomenon reflects error accumulation and is further validated experimentally, as discussed in detail in Section 3.4.
To solve this problem we propose the collaborative loss function as shown in Eq. 7, where $\hat{S}_{i}$, $s_{i}$ and $S_{i}$ are anomaly scores of $i^{t}h$ sample output by the conditional network, TSADM and LLM respectively. In Lemma 1, we prove that using the collaborative loss function is approximate to using the loss function in Eq. 6. This equivalence shows that Eq. 7 prevents error accumulation between the TSADM and LLM, as Eq.6 directly minimizes the difference between the collated anomaly scores of two samples and the ground truth. Additionally, we experimentally confirm that Eq. 7 effectively avoids error accumulation, as discussed in Section. 3.4.
$\begin{split}\mathcal{L}_{\beta}=-\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}\lambda_{1}&[(s_{i}-s_{j})(\hat{S}_{i}-\hat{S}_{j})]\\ &+\lambda_{2}[(S_{i}-S_{j})(\hat{S}_{i}-\hat{S}_{j})]\end{split}$ (7)
Lemma 1. When Assumption 1 - Assumption 3 hold, as the iteration step $T$ approaches infinity, minimizing Eq. 7 by stochastic gradient descent can be approximate to minimizing Eq. 6 by stochastic gradient descent with convergence rate $\mathbf{O}({T^{-\frac{1}{4}}})$.
The proof of Lemma 1 is given in Appendix. A.3.
Moreover, in Theorem 2, we prove that the optimal solution of Eq. 6 satisfies two key properties. Property 1 denotes that if the anomaly degree of time slot $r$ is more serious than the one of time slot $k$, the optimal collated anomaly score of time slot $r$ is bigger than that of time slot $k$. Property 2 denotes that if the difference in anomaly severity between time slots $r$ and $k$ is greater than the difference between time slots $r^{\prime}$ and $k^{\prime}$, the difference in optimal collated anomaly scores will accurately reflect this gap.
Theorem 2. When Assumption 1 - Assumption 3 hold, the optimal solution $\hat{S}^{*}$ of the loss function in Eq. 6 satisfies the following properties.
Property 1. $\forall r,\forall k$, if the ground truth $y_{r}>y_{k}$, then $\hat{S}^{*}_{r}>\hat{S}^{*}_{k}$, where $y_{r}$ and $y_{k}$ denotes the ground truth of the anomaly scores of $r^{th}$ and $k^{th}$ samples respectively, $\hat{S}^{*}_{r}$ and $\hat{S}^{*}_{k}$ are the ground truth of the anomaly scores of $r^{th}$ and $k^{th}$ samples respectively.
Property 2. $\forall r,\forall k,\forall r^{\prime},\forall k^{\prime}$, if the ground truth $y_{r}-y_{k}>y_{r^{\prime}}-y_{k^{\prime}}$, then $\hat{S}^{*}_{r}-\hat{S}^{*}_{k}>\hat{S}^{*}_{r^{\prime}}-\hat{S}^{*}_{k^{\prime}}$.
The proof of Theorem 2 is given in Appendix. A.4.
3 Experiment
Table 2: The average performance of CoLLaTe and baselines on four datasets.
Mustang MGAB Flight1 Flight2
Prec Rec F1 Prec Rec F1 Prec Rec F1 Prec Rec F1
GPT4 0.450 0.797 0.448 0.813 0.970 0.862 0.754 0.673 0.620 0.322 1.000 0.422
LLMAD 0.068 0.800 0.122 0.077 0.696 0.132 0.138 0.982 0.226 0.171 0.873 0.274
sigLLM - - - 0.063 0.589 0.107 - - - - - -
MSCRED 0.871 0.960 0.896 0.829 0.960 0.882 0.690 0.947 0.768 0.502 0.987 0.595
OmniAnom 0.812 0.968 0.878 0.761 0.995 0.822 0.619 1.000 0.721 0.579 1.000 0.637
TranAD 0.865 0.918 0.867 0.301 0.500 0.367 0.340 1.000 0.497 0.225 1.000 0.367
AnomalyTr 1.000 0.891 0.935 0.934 0.807 0.851 0.750 0.649 0.685 0.549 0.469 0.484
DCdetector 0.968 0.718 0.799 0.200 0.200 0.200 0.759 0.615 0.664 0.749 0.742 0.746
CoLLaTe 0.914 1.000 0.953 0.927 0.996 0.959 0.606 1.000 0.729 0.682 0.762 0.627
CoLLaTe 0.615 0.933 0.738 0.490 0.996 0.578 0.675 0.830 0.667 0.416 0.991 0.529
CoLLaTe 0.968 0.933 0.943 0.522 0.970 0.594 0.631 0.894 0.648 0.436 0.755 0.530
CoLLaTe 0.950 0.943 0.938 0.925 0.996 0.957 0.566 1.000 0.694 0.775 0.819 0.702
CoLLaTe 1.000 0.943 0.967 0.970 0.996 0.982 0.790 0.978 0.866 0.897 0.891 0.883
Table 3: The average performance of CoLLaTe and baselines on the unseen datasets.
Mustang MGAB Flight1 Flight2
Prec Rec F1 Prec Rec F1 Prec Rec F1 Prec Rec F1
MSCRED 0.719 0.984 0.804 0.671 0.963 0.765 0.734 0.956 0.765 0.426 0.983 0.521
OmniAnom 0.483 0.956 0.605 0.823 0.995 0.881 0.329 0.964 0.455 0.223 0.948 0.285
TranAD 0.308 0.501 0.354 0.321 0.600 0.395 0.188 0.600 0.282 0.090 0.400 0.147
AnomalyTr 0.200 0.200 0.200 0.976 0.667 0.765 0.391 0.251 0.305 0.592 0.458 0.491
DCdetector 0.250 0.250 0.250 0.325 0.399 0.353 0.408 0.295 0.325 0.750 0.747 0.748
CoLLaTe 0.950 0.943 0.938 0.916 0.996 0.948 0.796 1.000 0.877 0.897 0.891 0.883
We make extensive experiments on four datasets and make the following contributions:
  • CoLLaTe can achieve the best performance compared with SOTA LLM-based and TSADMs.
  • CoLLaTe can maintain good performance on unseen distributions.
  • We locate the hyperparameters important to CoLLaTe performance and analyze their effect.
  • All the modules in CoLLaTe are effective and contribute to its performance.
(a) Performance of CoLLaTe, TSADM and LLM
(b) The impact of patch length and learning rate
(c) The impact of $d$ and learning rate
(d) KL distance between the anomaly scores of the two models as iteration step grows up
Figure 3: (a) We use coordinates (Precision, Recall) to denote performance on a subset of flight dataset. We draw a scatter plot to show the distribution of LLM (GPT4), TSADM (CoLLaTe), and CoLLaTe; (b) The figure shows changes in F1 score as patch length in $D_{inter},D_{intra}$ and learning rate vary, where the learning rate ticks are the original value multiplied by 1000 times; (c) The figure shows changes in F1 score as $d$ and learning rate vary, where the learning rate ticks are the original value multiplied by 1000 times; (d) The figure shows that KL-divergence between the histogram of LLM anomaly score and aligned TSADM anomaly score decreases as the iteration step grows and compares the KL-divergence between LLM anomaly score and aligned TSADM anomaly score with the KL-divergence between LLM anomaly score and original TSADM anomaly score;
3.1 Experiment Setup
Baselines. We compare CoLLaTe with state-of-the-art (SOTA) anomaly detection methods, including DCdetector Yang et al., (2023), AnomalyTransformer Xu et al., (2022), TranAD Tuli et al., (2022), OmniAnomaly Su et al., (2019), and MSCRED Zhang et al., (2019). Additionally, we evaluate against SOTA LLM-based anomaly detection methods, including GPT-4, LLMAD Liu et al., 2024b , and sigLLM Alnegheimish et al., (2024). To better understand the contributions of each component in CoLLaTe, we conduct ablation studies by introducing several variations of the model. Specifically, CoLLaTe refers to the version of CoLLaTe without the alignment module. CoLLaTe replaces the collaborative loss with MSE. CoLLaTe modifies the weights $\lambda_{1}$ and $\lambda_{2}$ to $1$. Lastly, CoLLaTe removes the LLM and the conditional network entirely.
Datasets. To verify CoLLaTe can effectively embed the expertise knowledge, we choose three datasets for a complex system, which has domain-specific rules for anomaly detection and highly rely on expert knowledge. One is cluster-wide slowdown detection in cloud environment Amvrosiadis et al., (2018); Chen et al., 2024b and two are aircraft monitoring datasets collected from an aircraft manufacturing company that is one of the world’s top 150 enterprises. Additionally, we included a subtle mathematical-assumption-based anomaly detection dataset MGAB Thill et al., (2020), in which anomalies can not be recognized visually, but need complex mathematical derivation and verification. For more details, please refer to Appendix A.7.
Hyperparameters. We show the hyperparameter values in Appendix A.8.
Evaluation Metrics. We use three of the most popular evaluation metrics to measure the performance of CoLLaTe, as many works did Chen et al., 2022b ; Chen et al., (2021); Li et al., (2023): the recall, precision and F1 score.
3.2 Prediction Accuracy
Precision, recall, and F1 score are calculated for each subset, and the average results are presented in Tab. 2, where the best performance is highlighted in bolded, and if CoLLaTe achieves the best performance we underline the best performance among baselines. In the table, “prec" and “rec" denote precision and recall, while “OmniAnom" and “AnomalyTr" represent “OmniAnomaly" and “AnomalyTransformer", respectively. As sigLLM is specifically designed for univariate time series, it cannot be applied to the Mustang, Flight 1 and Flight 2 datasets, and therefore, its performance on these datasets is not included. As shown in Tab. 2, CoLLaTe achieves the highest F1 scores across all four datasets. In this table, GPT-4 represents the LLM method employing the set-up-pitch prompt, which is the LLM component of CoLLaTe. By comparing the performance of GPT-4 with other LLM-based anomaly detection methods, such as sigLLM and LLMAD, it becomes evident that the set-up-pitch prompt significantly enhances the anomaly detection accuracy of the LLM. In Fig. 3(a), we plot the precision and recall corresponding to the best F1 scores as coordinates to illustrate the performance distribution of CoLLaTe, GPT-4, and CoLLaTe. The figure reveals that both the LLM and the TSADM (CoLLaTe) tend to suffer from either low precision or low recall. In contrast, CoLLaTe demonstrates strong performance on both metrics. This observation confirms that CoLLaTe enables effective collaboration between the TSADM and the LLM, resulting in superior overall performance.
Since the datasets consist of different subsets and different subsets represent monitoring data for different cloud servers and aircraft components, they exhibit different distributions. To verify that LLM component in CoLLaTe can promote its performance on unseen distribution by leveraging general expertise knowledge, in Tab. 3, for each subset, we train the models on other subset and test the model on it. In this situation, CoLLaTe can maintain its strength over other methods, even for some baselines like TranAD equipping with meta-learning to improve its generalization. Thus, LLM component in CoLLaTe can effectively leverage the general expertise knowledge.
3.3 Hyperparameter Sensitivity
We investigate the impact of several factors on CoLLaTe’s performance, including $d$ in Eq.27, patch size for $D_{inter}$ and $D_{intra}$, learning rate, and $\hat{\lambda}_{1}$ and $\hat{\lambda}_{2}$ in Eq.3. Notably, when $\hat{\lambda}_{1}$ and $\hat{\lambda}_{2}$ vary within $[0.01,0.9]$, the change in CoLLaTe’s F1 score is within 0.01. Thus, our analysis primarily focuses on the effects of $d$, patch size, and learning rate. In Fig.3(b), we examine CoLLaTe’s F1 score as the patch size of $D_{inter}$ and $D_{intra}$ varies from 2 to 7, and the learning rate changes from 0.001 to 0.007. As shown in Fig.3(b), the F1 score improves as both the patch size and learning rate decrease. CoLLaTe performs better with smaller patch sizes because larger patch sizes that exceed the length of contextual anomalies lead to smaller $D_{inter}$. In such cases, since both contextual anomalies and point anomalies have small $D_{inter}$ and large $D_{intra}$, distinguishing between these anomaly types becomes difficult, thereby degrading CoLLaTe’s performance. In Fig.3(c), we evaluate CoLLaTe’s F1 score as $d$ ranges from 0.01 to 2 and the learning rate varies from 0.001 to 0.007. As illustrated in Fig.3(c), the F1 score increases as $d$ grows. Additionally, CoLLaTe’s performance becomes more stable across different learning rates as $d$ increases. Furthermore, smaller learning rates generally lead to better performance when $d$ is within the range $[0.01,2]$.
3.4 Effectiveness of Each Module
Ablation experiment. In CoLLaTe, we remove the Alignment module. In CoLLaTe, we change $\mathcal{L}(a,b)$ to MSE. In CoLLaTe, we change $\frac{D_{intra}}{D_{inter}+D_{intra}}$ and $\frac{D_{inter}}{D_{inter}+D_{intra}}$ to 1. We compare their performance with CoLLaTe in Tab. 2, where CoLLaTe outperforms all of them. This proves that each module in CoLLaTe contributes to its performance.
Effectiveness of the collaborative loss. CoLLaTe is the TSADM utilized within CoLLaTe. By comparing the performance of CoLLaTe, GPT-4, and CoLLaTe, it is evident that the collaboration between the TSADM CoLLaTe and the LLM-based model GPT-4 allows CoLLaTe to outperform both. This demonstrates that CoLLaTe effectively mitigates error accumulation during the collaboration between GPT-4 and CoLLaTe and coincides with the theoretical analysis in Theorem 1.
Effective of the design of $\mathcal{L}(a,b)$. Moreover, as mentioned earlier, CoLLaTe refers to the method that replaces $\mathcal{L}(a,b)$ in the collaborative loss function with MSE. On the MGAB dataset, the performance of CoLLaTe is worse than both GPT-4 and CoLLaTe. On other datasets, its performance is consistently worse than one of GPT-4 or CoLLaTe while being only slightly better than the other. This observation supports the analysis in Section 2.3, which indicates that using MSE as the formulation of $\mathcal{L}(a,b)$ leads to error accumulation between the LLM and the TSADM during the collaboration process.
Effectiveness of the adaptive hyperparameters. In CoLLaTe, the adaptive hyperparameters $\frac{D_{intra}}{D_{inter}+D_{intra}}$ and $\frac{D_{inter}}{D_{inter}+D_{intra}}$ are set to 1. When comparing CoLLaTe to GPT-4 and the TSADM CoLLaTe, we observe that on most datasets (MGAB, Flight 1, Flight 2), the performance of CoLLaTe falls between that of the LLM (GPT-4) and the TSADM (CoLLaTe). This is because, without effectively leveraging the complementary strengths of the LLM and the TSADM, the collaborative performance of CoLLaTe becomes a compromise, reflecting a balance between the better-performing and lower-performing models. This observation highlights the importance and effectiveness of the adaptive hyperparameters.
Effectiveness of the alignment module. The ablation experiment of CoLLaTe demonstrates that the alignment module contributes significantly to the performance of CoLLaTe. To verify that the alignment module effectively reduces the distribution distance between the aligned anomaly scores of the TSADM and the anomaly scores of the LLM, we compute the KL-divergence distance between the aligned scores of the TSADM and the anomaly score histogram of LLM as the iteration step grows. We present the results in Fig. 3(d). The figure also compares the KL-divergence distance between aligned TSADM anomaly scores and that of LLM with the KL-divergence distance between original TSADM anomaly scores and that of LLM. From Fig. 3(d), we observe that the distribution distance between the aligned TSADM anomaly scores and the LLM anomaly scores gradually decreases as the iteration steps increase. Eventually, it falls below the distribution distance observed between the original anomaly scores of the two models. This result verifies that the alignment module effectively aligns the distribution of anomaly scores between the TSADM and the LLM.
4 Related work
Anomaly detection methods generally fall into two categories: classical anomaly detection methods and neural network-based methods. Classical approaches primarily utilize statistical models Eskin, (2000); Wang et al., (2016), linear models Ma and Perkins, (2003); Amer et al., (2013); Hu et al., (2003), and proximity analysis Li et al., (2021); Izakian and Pedrycz, (2014) to identify anomalies. Nevertheless, these methods can not effectively leverage the expert knowledge.
Neural network-based methods, on the other hand, can be categorized into prediction-based methods Malhotra et al., (2015); Hundman et al., (2018); Zong et al., (2018); Chen et al., 2022a , reconstruction-based methods Chen et al., 2022b ; You et al., (2022); Jiang et al., (2022); Shen et al., (2021); Tian et al., (2019), and LLM-based methods Liu et al., 2024b ; Russell-Gilbert et al., (2024); Liu et al., 2024a . Prediction-based methods identify anomalies by comparing predicted values to ground truth, but they are vulnerable to historical inference errors Wang et al., (2024). Reconstruction-based methods address this limitation. However, they cannot leverage expert knowledge in specific application scenarios without model adaptation. These methods identify anomalies by learning normal patterns from training data and flagging data samples that deviate from these patterns. Yet, in scenarios where it is impossible to observe or collect all normal patterns (e.g., airplane monitoring), reconstruction-based methods are ineffective. In contrast, LLM-based methods introduce expert knowledge effectively through techniques like retrieval-augmented generation (RAG) Lewis et al., (2020) and chain of thought (COT) Wei et al., (2022) without requiring model modifications. However, as natural language-based models, LLMs are insensitive to value fluctuations and normal pattern extraction. None of these existing methods fully leverage the complementary strengths of LLM-based approaches and TSADM.
5 Conclusion
Given the complementary strengths of LLMs and TSADM, where LLMs can conveniently incorporate expert knowledge, while TSADM are more sensitive to value fluctuations and normal pattern extraction, we propose a novel approach inspired by the human nervous system to facilitate their collaboration. We begin by identifying the key challenges in facilitating this collaboration and then introduce a collaborative framework, called CoLLaTe, to address them. There are two distinctive characteristics of CoLLaTe: an alignment module aligning the expression domain of LLM and TSADM, and a collaborative loss function reducing the error accumulation. We have made solid mathematical proof and extensive experiments to verify the effectiveness of the proposed methods.
Appendix A Appendix
A.1 Proof of Transformation
[…]
In this part, we prove formation in Eq. 8 can be transformed to Eq. 3, when let $N$ approach infinite. When the $N$ approaches infinite, the optimization objective in Eq. 8 can be transformed to $\min.\lim_{N\to\infty}-\sum_{i=1}^{N}\frac{c_{i}}{\sum_{j=1}^{N}c_{j}}\log f(\frac{ib}{N})$. For each $c_{i}$, there will be two situations. One is that there is no $\mathcal{M}(s_{j})=\frac{ib}{N},j\in[1,n]$, then $c_{i}=0$ and the item including $c_{i}$ can be omitted. Another one is that there are some $\mathcal{M}(s_{j})=\frac{ib}{N},j\in[1,n]$, then $c_{i}$ is equal to the number of times that $\frac{ib}{N}$ appears in $\{\mathcal{M}(s_{j})|s_{j}\in\{s_{j}\}_{j=1}^{n}\}$. In this situation, the original item $\frac{c_{i}}{\sum_{j=1}^{N}c_{j}}\log f(\frac{ib}{N})$ can be transformed to $\frac{c_{i}}{n}\log f(\mathcal{M}(s_{j}))$, since $\sum_{j=1}^{N}c_{j}=n$. We cluster the mapping scores with the same value as a group, from which we can obtain $M$ groups with values of $\{\mathcal{M}(s_{\hat{j}})\}_{\hat{j}=1}^{M}$. The size of $\hat{j}^{th}$ group is exactly $c_{i}$, where $ib/N=\mathcal{M}(s_{\hat{j}})$. We denote the $c_{i}$ as $c_{\hat{j}}$. The $\{c_{\hat{j}}\}_{\hat{j}=1}^{M}$ consist of the non-zero part of $\{c_{i}\}_{i=1}^{N}$. Then the original formation can be transformed to $\min.-\sum_{\hat{j}=1}^{M}\frac{c_{\hat{j}}}{n}\log f(\mathcal{M}(s_{\hat{j}}))$, which is exactly equal to traversing all the $\mathcal{M}(s_{i}),s_{i}\in\{s_{i}\}_{i=1}^{n}$, sum them up as $\sum_{i=1}^{n}\log(f(\mathcal{M}(s_{i})))$, and combine the items with same value of mapping scores.
A.2 Proof of Theorem 1
When $\mathcal{L}(a,b)=(a-b)^{2}$, we can transform Eq. 5 to Eq. 9, where $\lambda_{1}=\frac{D_{intra}}{D_{intra}+D_{inter}}$ and $\lambda_{2}=\frac{D_{inter}}{D_{intra}+D_{inter}}$.
$\lambda_{1}(y+\epsilon_{s}-\hat{S}(\mathcal{P}))^{2}+\lambda_{2}(y+\epsilon_{S}-\hat{S}(\mathcal{P}))^{2}$ (9)
According to KKT condition Boyd and Vandenberghe, (2004), the optimal solution of $\mathcal{L}_{\beta}$ should satisfy Eq. 10.
$0\displaystyle 2\lambda_{1}(y+\epsilon_{s}-\hat{S}^{*}(\mathcal{P}))\frac{\partial\hat{S}^{*}(\mathcal{P})}{\partial\mathcal{P}}+2\lambda_{2}(y+\epsilon_{S}-\hat{S}^{*}(\mathcal{P}))\frac{\partial\hat{S}^{*}(\mathcal{P})}{\partial\mathcal{P}}=$ 0 (10)
As shown in Appendix A.9, the gradient of architecture of conditional network can not be equal to 0. Thus, $02\lambda_{1}(y+\epsilon_{s}-\hat{S}^{*}(\mathcal{P}))+2\lambda_{2}(y+\epsilon_{S}-\hat{S}^{*}(\mathcal{P}))=$ 0. Then, $\hat{S}^{*}(\mathcal{P})=y+\lambda_{1}\epsilon_{s}+\lambda_{2}\epsilon_{S}$. Thus, $\mathbb{E}[(\hat{S}^{*}(\mathcal{P})-y)^{2}]=\mathbb{E}[(\lambda_{1}\epsilon_{s}+\lambda_{2}\epsilon_{S})^{2}]$. Since $f(x)=x^{2}$ is a convex function, according to Jensen’s Inequality, $\mathbb{E}[(\lambda_{1}\epsilon_{s}+\lambda_{2}\epsilon_{S})^{2}]\geq[\mathbb{E}(\lambda_{1}\epsilon_{s}+\lambda_{2}\epsilon_{S})]^{2}$. Since $\epsilon_{1}\sim\mathcal{D}_{s}(\mu_{s},\sigma_{s}),\epsilon_{2}\sim\mathcal{D}_{S}(\mu_{S},\sigma_{S})$, we have $[\mathbb{E}(\lambda_{1}\epsilon_{s}+\lambda_{2}\epsilon_{S})]^{2}=(\lambda_{1}\mu_{1}+\lambda_{2}\mu_{2})^{2}$. Thus, $\mathbb{E}[(\hat{S}^{*}(\mathcal{P})-y)^{2}]\geq(\lambda_{1}\mu_{1}+\lambda_{2}\mu_{2})^{2}$
A.3 Proof of Lemma 1
When using stochastic gradient descent algorithm, the gradient of Eq. 5 and Eq. 6 is shown in Eq. 11 and Eq. 12 respectively, where $\mathcal{B}$ is the set of sample in batch and $|\mathcal{B}|$ is the size of set $\{(r,t)|r\in\mathcal{B},t\in\mathcal{B}\}$. Let $\mathbf{g}_{i,j}^{*}$ denote the item in Eq. 12, whose index is $(i,j)$. Let $\mathbf{g}_{i,j}$ denote the item in Eq. 11, whose index is $(i,j)$. $\mathbf{g}_{i,j}$ can be transformed to $\mathbf{g}_{i,j}^{*}+[\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j})](\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}$. Thus, the gradient of Eq. 5 can be deemed as the gradient of Eq. 6 add some random noise scaled by $(\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}$ for each sample. Let $\hat{S}^{{\dagger}}$ denote the optimal solution of Eq. 6. Thus, $\mathcal{L}_{\beta}^{*}(\hat{S})\geq\mathcal{L}_{\beta}^{*}(\hat{S}^{{\dagger}}),\forall\hat{S}$, and it satisfies Assumption 5.1 in Bu et al., (2024).
$\begin{split}\mathbf{g}=-\frac{1}{|\mathcal{B}|}\sum_{(i,j)\in\{(r,t)|r\in\mathcal{B},t\in\mathcal{B}\}}\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j}+y_{i}-y_{j})(\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}\\ +\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j}+y_{i}-y_{j})(\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}\end{split}$ (11)
$\begin{split}\mathbf{g}^{*}=-\frac{1}{|\mathcal{B}|}\sum_{(i,j)\in\{(r,t)|r\in\mathcal{B},t\in\mathcal{B}\}}\lambda_{1}(y_{i}-y_{j})(\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}\\ +\lambda_{2}(y_{i}-y_{j})(\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}\end{split}$ (12)
Since $y\in[0,1],\hat{S}\in[0,1]$, $|y_{i}-y_{j}|\leq 1,\forall(i,j)$ and $|\hat{S}_{i}-\hat{S}_{j}|\leq 1,\forall(i,j)$. Thus, $|\mathbf{g}_{i,j}^{*}|=|(y_{i}-y_{j})(\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}|\leq|\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}|$. Furthermore, according to Appendix. A.9, conditional network is a multi-layer perceptron with sigmoid as its activation function. Let $\mathbf{\sigma}(x)$ denote sigmoid function. Since $\frac{\partial\mathbf{\sigma}(x)}{x}=\mathbf{\sigma}(x)(1-\mathbf{\sigma}(x))$ and $10<\mathbf{\sigma}(x)<$ 1, $|\frac{\partial(\hat{S})}{\partial\mathcal{P}}|<1$. Thus, $|\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}|\leq 1$. Thus, $|\mathbf{g}_{i,j}^{*}|\leq 1$. Thus, $|\mathbf{g}^{*}|\leq\frac{1}{|\mathcal{B}|}\sum_{i,j}|\mathbf{g}_{i,j}^{*}|=1$. Since $\mathcal{L}_{\beta}^{*}$ is $L$-Lipschitz continuous, it satisfies the inequality in Eq. 13.
$\begin{split}|\mathcal{L}_{\beta}^{*}(\mathcal{P}_{1})-(\mathcal{L}_{\beta}^{*}(\mathcal{P}_{2})+{\mathbf{g}^{*}}^{T}(\mathcal{P}_{1}-\mathcal{P}_{2}))|=\\ |[\mathcal{L}_{\beta}^{*}(\mathcal{P}_{1})-\mathcal{L}_{\beta}^{*}(\mathcal{P}_{2})]+{\mathbf{g}^{*}}^{T}(\mathcal{P}_{1}-\mathcal{P}_{2})|\leq\\ L|\mathcal{P}_{1}-\mathcal{P}_{2}|+|\mathbf{g}_{i,j}^{*}|\cdot|\mathcal{P}_{1}-\mathcal{P}_{2}|\leq\\ (L+1)\cdot|\mathcal{P}_{1}-\mathcal{P}_{2}|\end{split}$ (13)
Since $L\geq 0$, $L+1>0$, our model satisfies Assumption 5.2 in Bu et al., (2024).
We can transform $\mathbb{E}(\mathbf{g}_{i,j}^{*}-\mathbf{g}_{i,j})$ step by step, as shown in Eq. 14.
$\begin{split}\mathbb{E}(\mathbf{g}_{i,j}^{*}-\mathbf{g}_{i,j})=\mathbb{E}[(\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j}))(\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}]\\ =\mathbb{E}[(\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j}))](\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}\end{split}$ (14)
Since $\mathbb{E}(\epsilon_{s,i})=\mathbb{E}(\epsilon_{s,j}),\mathbb{E}(\epsilon_{S,i})=\mathbb{E}(\epsilon_{S,j})$, we can obtain $\mathbb{E}[\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j})]=0$. Thus, $\mathbb{E}(\mathbf{g}_{i,j}^{*}-\mathbf{g}_{i,j})=0$. $\mathbb{E}|\mathbf{g}_{i,j}^{*}-\mathbf{g}_{i,j}|^{2}$ can be calculated as shown in Eq. 15.
$\begin{split}\mathbb{E}|\mathbf{g}_{i,j}^{*}-\mathbf{g}_{i,j}|^{2}=(\hat{S}_{i}-\hat{S}_{j})^{2}(\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}})^{2}\mathbb{E}[(\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j}))^{2}]\end{split}$ (15)
As proven above, $|\hat{S}_{i}-\hat{S}_{j}|\leq 1$, thus $(\hat{S}_{i}-\hat{S}_{j})^{2}\leq 1$. Similarly $\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}}\leq 1$, we can obtain $(\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}})^{2}\leq 1$. Thus, we can obtain the inequality in Eq. 16.
$\begin{split}(\hat{S}_{i}-\hat{S}_{j})^{2}(\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}})^{2}\mathbb{E}[(\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j}))^{2}]\leq\\ \mathbb{E}[(\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j}))^{2}]\end{split}$ (16)
Since $\epsilon_{s,i}\sim\mathcal{D}(\mu_{s},\sigma_{s}),\epsilon_{s,j}\sim\mathcal{D}(\mu_{s},\sigma_{s})$ independently, $\mathbf{E}[(\epsilon_{s,i}-\epsilon{s,j})^{2}]=2\sigma_{s}^{2}$. Similarly, $\mathbf{E}[(\epsilon_{S,i}-\epsilon_{S,j})^{2}]=2\sigma_{S}^{2}$. After simplifying its formation, we obtain $\mathbb{E}[(\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j}))^{2}]=2\lambda_{1}^{2}\sigma_{s}^{2}+2\lambda_{2}^{2}\sigma_{S}^{2}$. Thus, $\mathbb{E}|\mathbf{g}_{i,j}^{*}-\mathbf{g}_{i,j}|^{2}\leq 2\lambda_{1}^{2}\sigma_{s}^{2}+2\lambda_{2}^{2}\sigma_{S}^{2}$. Then, it satisfies Assumption 5.3 in Bu et al., (2024).
Therefore, $\mathcal{L}_{\beta}^{*}$ satisfies Assumption 5.1-Assumption 5.3 in Bu et al., (2024). As proven before, minimizing $\mathcal{L}_{\beta}$ by stochastic gradient descent can be deemed as adding a random noise to the gradient of each sample during the stochastic gradient descent process of $\mathcal{L}_{\beta}^{*}$. Since Assumption 5.1-Assumption 5.3 in Bu et al., (2024) hold, we can use Theorem 4 in Bu et al., (2024) to prove that after adding these random noise, $\mathcal{L}_{\beta}$ can also converge as quickly as $\mathcal{L}_{\beta}^{*}$ by $\mathbf{O}(T^{-\frac{1}{4}})$. Besides, we have the equation as shown in Eq. 17.
$\mathbb{E}(\mathbf{g}_{i,j})=\mathbb{E}(\mathbf{g}_{i,j}^{*})+\mathbb{E}([\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j})](\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}})$ (17)
Since $\mathbb{E}[\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j})]=0$, we can obtain $\mathbb{E}([\lambda_{1}(\epsilon_{s,i}-\epsilon_{s,j})+\lambda_{2}(\epsilon_{S,i}-\epsilon_{S,j})](\hat{S}_{i}-\hat{S}_{j})\frac{\partial(\hat{S}_{i}-\hat{S}_{j})}{\partial\mathcal{P}})=0$. Thus, $\mathbb{E}(\mathbf{g}_{i,j})=\mathbb{E}(\mathbf{g}_{i,j}^{*})$. Since $\mathcal{L}_{\beta}$ can converge as proven above, its gradient $\mathbb{E}(\mathbf{g})$ should converge to 00 as iteration step grows up. Thus, as iteration step grows up, $\mathbb{E}(\mathbf{g}_{i,j})$ approaches 0, i.e., $\mathcal{P}$ approaches the one that makes $\mathbb{E}(\mathbf{g}_{i,j}^{*})=0$. Since $\mathcal{L}_{\beta}^{*}$ is convex, as long as $\mathbb{E}(\mathbf{g}_{i,j}^{*})=0$, $\mathcal{P}$ is the optimal solution of $\mathcal{L}_{\beta}^{*}$. Thus, by using $\mathcal{L}_{\beta}$, $\mathcal{P}$ can converge to the optimal solution of $\mathcal{L}_{\beta}^{*}$ with similar converge rate $\mathbf{O}(T^{-\frac{1}{4}})$, as iteration step approaches infinite.
A.4 Proof of Theorem 2
Since when Assumption 1 - Assumption 4 hold, Lemma 1 is valid. Thus, using stochastic gradient descent (SGD) to minimize $\mathcal{L}_{\beta}$ can be approximate to using stochastic gradient descent to minimize $\mathcal{L}_{\beta}^{*}$, when iteration step approach infinity. Thus, we prove that optimal solution of $\mathcal{L}_{\beta}^{*}$ has the mentioned properties.
Proof of property 2. Let $\hat{S}^{{\dagger}}$ denote the optimal solution of $\mathcal{L}_{\beta}^{*}$. We use the proof of contradiction to prove it. We firstly assume if there are $r,k,r^{\prime},k^{\prime}\in[1,n]$ that satisfies Eq. 18-Eq. 19. By enlarging $\hat{S}^{{\dagger}}_{r}$ by $\Delta t$ and reducing $\hat{S}^{{\dagger}}_{r^{\prime}}$ by $\Delta t$, we can further decrease the value of $\mathcal{L}_{\beta}^{*}$, which is contradict to the assumption that $\hat{S}^{{\dagger}}$ is optimal. Thus, $\forall r,k,r^{\prime},k^{\prime}$, if $y_{r}-y_{k}>y_{r^{\prime}}-y_{k^{\prime}}$, then $\hat{S}^{{\dagger}}_{r}-\hat{S}^{{\dagger}}_{k}>\hat{S}^{{\dagger}}_{r^{\prime}}-\hat{S}^{{\dagger}}_{k^{\prime}}$.
Here is the detailed derivation. Without loss of generality, we can assume $y_{r}>y_{r^{\prime}}$, because if $y_{r}<y_{r^{\prime}}$, given $y_{r}-y_{k}<y_{r^{\prime}}-y_{k^{\prime}}$, $y_{k}<y_{k^{\prime}}$. In this case, by multiplying minus one to both side of the inequality Eq. 18-Eq. 19, we can obtain a set of inequality that $y_{k^{\prime}}-y_{r^{\prime}}>y_{k}-y_{r},\hat{S}^{{\dagger}}_{k^{\prime}}-\hat{S}^{{\dagger}}_{r^{\prime}}<\hat{S}^{{\dagger}}_{k}-\hat{S}^{{\dagger}}_{r}$ with the condition $y_{k^{\prime}}>y_{k}$, which is equal to assuming $y_{r}>y_{r^{\prime}}$ for Eq. 18-Eq. 19.
$y_{r}-y_{k}>y_{r^{\prime}}-y_{k^{\prime}}$ (18)
$\hat{S}^{{\dagger}}_{r}-\hat{S}^{{\dagger}}_{k}<\hat{S}^{{\dagger}}_{r^{\prime}}-\hat{S}^{{\dagger}}_{k^{\prime}}$ (19)
Let $\hat{S}^{{\dagger}}_{r}$ grows to $\hat{S}^{{\dagger}}_{r}+\Delta t$ and $\hat{S}^{{\dagger}}_{r^{\prime}}$ decreases to $\hat{S}^{{\dagger}}_{r^{\prime}}-\Delta t$, where $\Delta t>0$. Let $\mathcal{L}_{1}$ denote the value of $\mathcal{L}_{\beta}^{*}$ before changing $\hat{S}^{{\dagger}}_{r}$ and $\hat{S}^{{\dagger}}_{r^{\prime}}$ and $\mathcal{L}_{2}$ denote the value of $\mathcal{L}_{\beta}^{*}$ after changing $\hat{S}^{{\dagger}}_{r}$ and $\hat{S}^{{\dagger}}_{r^{\prime}}$. We compute the difference between $\mathcal{L}_{2}$ and $\mathcal{L}_{1}$ in Eq. 20. Furthermore, we can simplify it to Eq. 21. Given $y_{r}>y_{[}r^{\prime}]$, $\mathcal{L}_{1}-\mathcal{L}_{2}>0$, which means $\mathcal{L}_{2}$ is smaller than $\mathcal{L}_{1}$. However, since $\hat{S}^{{\dagger}}$ is the optimal solution of $\mathcal{L}_{\beta}^{*}$, there should be no value set of $\hat{S}$ that can make $\mathcal{L}_{\beta}^{*}$ smaller than $\mathcal{L}_{1}$. Thus, the conclusion $\mathcal{L}_{2}<\mathcal{L}_{1}$ is contradict to the assumption that $\hat{S}^{{\dagger}}$ is optimal. Thus, $\forall r,k,r^{\prime},k^{\prime}$, if $y_{r}-y_{k}>y_{r^{\prime}}-y_{k^{\prime}}$, then $\hat{S}^{{\dagger}}_{r}-\hat{S}^{{\dagger}}_{k}>\hat{S}^{{\dagger}}_{r^{\prime}}-\hat{S}^{{\dagger}}_{k^{\prime}}$.
$\begin{split}\mathcal{L}_{1}-\mathcal{L}_{2}=\sum_{i=1,i\neq r}^{n}(y_{i}-y_{r^{\prime}})\Delta t+\sum_{j=1,j\neq r}(y_{r^{\prime}}-y_{j})(-\Delta t)+(y_{r}-y_{r^{\prime}})(2\Delta t)+(y_{r^{\prime}}-y_{r})(-2\Delta t)\\ +\sum_{i=1,i\neq r^{\prime}}^{n}(y_{i}-y_{r})(-\Delta t)+\sum_{j=1,j\neq r^{\prime}}^{n}(y_{r}-y_{j})(\Delta t)\end{split}$ (20)
$\begin{split}\mathcal{L}_{1}-\mathcal{L}_{2}=2n\Delta t(y_{r}-y_{r^{\prime}})\end{split}$ (21)
Proof of Property 1. Let $k=k^{\prime}$. Then, according to property 2, $\forall r,k,r^{\prime}$, if $y_{r}-y_{k}>y_{r^{\prime}}-y_{k}$, $\hat{S}^{{\dagger}}_{r}-\hat{S}^{{\dagger}}_{k}>\hat{S}^{{\dagger}}_{r^{\prime}}-\hat{S}^{{\dagger}}_{k}$. The inequality mentioned above can further reduced to $\forall r,r^{\prime}$, if $y_{r}>y_{r^{\prime}}$, $\hat{S}^{{\dagger}}_{r}>\hat{S}^{{\dagger}}_{r^{\prime}}$.
A.5 Train and Inference
During the training process, we use $\mathcal{L}_{a}+\mathcal{L}_{\beta}$ as the loss function. During the inference process, we use the collated anomaly score output by the conditional network as the anomaly score.
A.6 Anomaly Types
We use different binary classification standards to categorize the anomalies in Flight 1 and Flight 2. Among them, the most common binary standard is point anomaly and contextual anomaly, which is illustrated in  Lai et al., (2024). Besides, we also introduce two other kinds of binary classification standards for aircraft monitoring datasets to explore the complementary performance between LLM and TSADM. According to the degree of reliance on expert knowledge for anomaly detection, we can divide anomalies into background anomalies and data pattern anomalies. The background anomalies highly rely on expertise knowledge, while the data pattern anomaly can be identified only by TSADMs. As shown in Fig. 4(a), the valid range of the dependent variable is decided by the independent variable, which should satisfy the laws of physics between different parameters. It is usually hard for a neural network to clearly identify the valid range of the dependent variable, given the independent variable. Thus, in this case, it relies on expert knowledge and belongs to background anomaly. In contrast, in Fig. 4(b), the redundant channel mechanism usually requires that most of the redundant channels can reach a consensus. Thus, different redundant channels should always be similar to each other in normal status. In this situation, the neural network can identify the similarity between different channels without expert knowledge. Thus, it belongs to data pattern anomaly. According to the similarity between anomaly and surrounding time series, we can divide anomalies into gradual anomalies and mutant anomalies. As shown in Fig. 4(c), the anomalies are quite different from surrounding normal time series and they should belong to mutant anomalies. In Fig. 4(d), the anomalies are similar to the surrounding normal time series and they should belong to the gradual anomaly.
(a) The background anomaly
(b) The data pattern anomaly
(c) The mutant anomaly
(d) The gradual anomaly
Figure 4: These figures show different anomaly types
A.7 Datasets
In our experiments, we used four datasets from different scenarios: cloud service monitoring, aircraft monitoring, and mathematical assumption-based anomaly datasets. We introduce these datasets in the following. The used datasets of MGAB and Mustang are included in "Supplementary Material". As for Flight 1 and Flight 2, since they refers to business secret, we could not disclose them.
  • MGAB.The dataset comprises a synthetic Mackey-Glass time series characterized by non-trivial anomalies. Mackey-Glass time series are recognized for their propensity to exhibit chaotic behavior under specific conditions. The MGAB dataset encompasses 10 Megabytes (MG) of time series data, with each series extending to a length of 105 time points. Within each time series, 10 anomalies are meticulously introduced using a procedure outlined subsequently. Unlike other synthetic benchmarks, it poses a considerable challenge for the human eye to discern the artificially introduced anomalies from the inherent chaotic behavior. A segment of a time series featuring 3 anomalies is illustrated in the graph above. Due to the limit of computation resources when using LLM, we only use 10000 time slots in our experiments. Since there are no point anomalies in the original dataset, but our model utilizes the complementary performance of LLM and TSADM on point anomaly and contextual anomaly, we insert some point anomalies into the original dataset and release it in "Supplementary Material".
  • Mustang. The Mustang dataset chronicles the task duration of operations over a span of five years. The authors in Chen et al., 2024b have conducted preprocessing on the raw dataset, where they count the histogram of the task duration time distribution at each time slot. Subsequently, they manually identified and labeled the slowdown anomalies. We include used dataset in "Supplementary Material".
  • Flight 1 & 2. These datasets are collected from an airline company, which is one of the world’s top 150 enterprises. Different data sets are monitoring data of different flight control system components. The datasets use redundant channels Osder, (1999) to ensure the robustness of flight systems and refer to sophisticated collaborative relationships between different parameters, which introduces complex rules to judge the availability of the flight control system and is hard to learn only by neural networks. Besides, we show the ratios of normality and different kinds of anomaly in Fig. 5(a), where the anomaly classification is defined in Appendix. A.6.
A.8 Experiment details
The four datasets used in our experiments consist of multiple subsets. For each subset, we use 40% as the training set, 10% as the validation set, and the remaining 50% as the testing set. We illustrate the meaning of each hyperparameter in Tab. 4 and list the value of the hyperparameter in Tab. 5. Since the datasets used in our experiments consist of several subsets. In some datasets, we set different hyperparameter values for different subsets and we list the values used in the Tab. 5. In these situations, we use grid search to find the optimal combinations of these hyperparameters for each subset.
Table 4: The meaning of each hyperparameter.
Hyperparameter Name Meaning
winLen The length of sliding window
moduleNum The number of TSADM attention layers
batchSize The size of batch
trlr The learning rate of TSADM
colr The learning rate during collaboration
patchSize The size of patch in computin $D_{inter}$ and $D_{intra}$
kLen The length of convolution kernel in TSADM
$d$ The square root in Eq. 27
$\hat{\lambda}$ The hyperparameter weight in $\mathcal{L}_{a}$
Table 5: The value of each hyperparameter.
Mustang MGAB Fligh 1 Flight 2
Name Value Name Value Name Value Name Value
winLen 4 winLen 5 winLen 2 winLen 2
moduleNum 6 moduleNum 10 moduleNum 6 moduleNum 10
batchSize 100 batchSize 100 batchSize 100 batchSize 100
trlr 0.001 trlr 0.01 trlr 0.01 trlr 0.01
colr 0.001,0.0001 colr 0.01 & 0.001 colr 0.008 colr 0.01,0.001
patchSize 2 patchSize 2 patchSize 2 patchSize 2
kLen 2 kLen 2 kLen 2 kLen 2
d 0.7,1.2 d 0,0.5,0.7,1,1.5 d 0,1.7,2 d 0.01,0.5,0.9,1.7,2
$\hat{\lambda}$ 0.1,1 $\hat{\lambda}$ 1 $\hat{\lambda}$ 1,2,5 $\hat{\lambda}$ 1,5
A.9 Architecture of Our Conditional Network
The conditional network uses the data representation obtained from a TSADM as the condition and leverages the aligned anomaly score of the TSADM and anomaly score of LLM to output the collated anomaly score. As shown in Eq. 22, we firstly concat data representation, aligned anomaly score of TSADM and anomaly score of LLM as many conditional networks do Rakelly et al., (2018). After that, we use multilayer perceptrons to output collated anomaly score as shown in Eq. 23, where $W_{1}$ and $W_{2}$ are trainable parameters, $\sigma_{1}$ and $\sigma_{2}$ are activation functions, such as Leaky ReLU.
$\bar{S}=\operatorname{Concat}(S,s,R)$ (22)
$\hat{S}=\sigma_{2}(W_{2}\sigma_{1}(W_{1}\cdot\bar{S}))$ (23)
A.10 Architecture of TSADM
The TSADM consists of 6-layer anomaly attention. Each anomaly attention layer is described in Eq. 24, where $G[i,j]=1-\exp^{-\frac{(i-j)^{2}}{\sigma^{2}}}$, $\sigma$ is a learnable parameter and * denotes an element-wise multiplication. This attention mechanism is built on the observation that anomaly could not establish decentralized connection with the whole time series Xu et al., (2022) and embeds moderate modification compared with the original one. To promote the anomaly attention capacity of dealing with compound periodic time series, different anomaly attention layers are organized by skimming attention mechanism Chen et al., 2024b .
$\mathcal{A}=q_{d}k_{d}^{T},$ (24)
$\tilde{\mathcal{A}}=\operatorname{softmax}(\mathcal{A}*G),$ (25)
$\tilde{x}_{l}[:,d]=\tilde{\mathcal{A}}v_{d}.$ (26)
A.11 Complementary Performance
In CoLLaTe, we mainly utilize the complementary performance of LLM and TSADM on point anomaly and contextual anomaly. Actually, there are other complementary performances between the TSADM and LLM by using another binary classification standard, and we explore them in this section. We introduce two kinds of binary classification standards: the anomalies that need expert knowledge to detect and the anomalies that can be detected only by data patterns, gradual anomalies, and mutual anomalies. These anomalies are illustrated in detail in Appendix. A.6. We divide Flight 2 into pieces and gather the pieces only obtain only one kind of anomaly. We test GPT4 and CoLLaTe on them and compute their average F1 score, which is shown in Fig. 5(c). In Fig. 5(c), we can see that LLM always has better performance on one kind of anomaly and has poorer performance on another in each binary anomaly classification.
(a) The distribution of Flight datasets
(b) GPT4 performance for different distances from examples
(c) The complementary strength of LLM and TSADM
(d) Validity of Assumption 3
Figure 5: (a) shows the distribution of flight datasets. (b) shows the anomaly detection performance of GPT4 when the distance between the designated time slot for anomaly detection and the time slot of the examples increases. (c) The figure shows the F1 score of LLM (GPT4) and TSADM (CoLLaTe) when using different anomaly binary classification criteria. (d) We verify that the mould length of $|\mathcal{L}_{\beta}^{*}(\theta_{1})-\mathcal{L}_{\beta}^{*}(\theta_{2})|_{2}$ is less than $L|\theta_{1}-\theta_{2}|_{2}$ and verifies the validity of assumption 3.
A.12 Set-up-pitch Prompt
Numerous studies have shown that incorporating similar historical examples into prompts can significantly enhance anomaly detection performance Liu et al., 2024b . Based on this insight, our prompt is structured into four components: expertise supplementation, task description, input data, and examples. The expertise supplementation component provides the contextual meanings of the various dimensions of the input data within the application domain, along with relevant domain-specific knowledge in professional document, such as typical value ranges and interrelationships between dimensions. The task description component directs the large language model (LLM) to determine whether a given time slot is anomalous and specifies the required output format. The examples component includes both a positive and a negative example to guide the model’s understanding. The detailed prompt design is provided in the following.
While recent works utilize examples from a fixed dataset, many application scenarios exhibit dynamic patterns over time. For instance, in cloud server monitoring, the cloud environment evolves with service updates, deployments, and revocations Chen et al., 2024a ; Ma et al., (2021). Similarly, in aircraft monitoring, normal patterns shift with changes in flight conditions and aircraft states. As illustrated in Fig. 5(b), our experiments with aircraft monitoring data reveal that the performance of GPT-4 degrades as the temporal distance between the anomaly detection time slot and the example time slot increases. To address this, we leverage a TSADM for anomaly detection to periodically label data flow and update the example collection with manual assistance, ensuring that the examples remain relevant and effective over time.
We list the concrete prompts for different datasets below. Since the prompt for aircraft monitoring datasets (Flight 1 and Flight 2) refers to the trade secrets of the airline company, we do not list them below. Due to the randomness of LLM, we ran multiple times and picked the one with the best performance to collaborate with the TSADM.
Prompt of MGAB
Expertise Supplement
: The input data is a $10000*1$ time series. The time series is generated by adding the value x, which satisfied $\frac{dx}{dt}=0.25*\frac{x(t-18)}{1+x(t-18)^{10}}-0.1*x(t)$, and some noises within range of $[-0.01,0.01]$, where $x(t)$ represents the value of time series at $t^{th}$ time slot. There are some anomalies inserted into the time series, which do not obey mentioned rules. The anomalies are inserted by repeating some future segments of the time series to the present position. [Professional document can be inserted into this part]
Input data: {data}
Task description: Given the input data, please pick out the anomalies along the time series and output the possibility that the $i^{th}$ time slot is anomalous. Please only output the corresponding possibility within the range of $[0,1]$. ($i$ is an iterable index).
Examples:
Example 1: Given the input data, please pick out the anomalies along the time series and output the possibility that the $i^{th}$ time slot is anomalous. Please only output the corresponding possibility within the range of $[0,1]$. Output: 0.
Example 2: Given the input data, please pick out the anomalies along the time series and output the possibility that the $i^{th}$ time slot is anomalous. Please only output the corresponding possibility within the range of $[0,1]$. Output: 1.
Prompt of Mustang
Expertise Supplement
: The input data is a matrix with size T * 17, which represents the task duration time distribution for T time slots in the cloud center. For each time slot, there are 17 dimensions. The values for these dimensions respectively represent the ratio of tasks whose duration time belongs to [0,5), [5,10), [10, 20), [20,30), [30, 40), [40, 70), [70, 110), [110, 150), [150, 190), [190, 230), [230, 280), [280, 330), [330, 380), [380, 430), [430, 900), [900, 1200), [1200, 1900). [Professional document can be inserted into this part]
Input data: {data}.
Task description: Given a sliding window with size $50*17$, please judge if there are many tasks slowdown at a specific time slot, compared with other slots in a given window. Please output a float number ranging from 0 to 1 to represent the probability that there are many tasks slowdown at the specific time slot.
Examples:
Example 1: Given the input data, please judge if there are many tasks that slow down at time slot $i$ and output a float number ranging from 0 to 1 to represent the probability of it. Response: 0.
Example 2: Given the input data, please judge if there are many tasks that slow down at time slot $i$ and output a float number ranging from 0 to 1 to represent the probability of it. Response: 1
A.13 Min-Max normalization to TSADM anomaly scores
we firstly scale anomaly scores from the TSADM by Eq. 27, where $d$ is a hyperparameter, $\dot{s}_{max}$ and $\dot{s}_{min}$ are the maximum and minimum values of these anomaly scores.
$s=\frac{\dot{s}}{\sqrt[d]{\dot{s}_{max}-\dot{s}_{min}}}$ (27)
A.14 Model Efficiency
Since we call the GPT-4 API, which will force the program to sleep for a while after sending several requests consecutively. Therefore, using a locally deployable LLM would significantly reduce the time compared to the times we have listed.
When using a K80 to train and inference CoLLaTe (LLM part calls API), the average time to obtain LLM anomaly scores is 1.27 seconds per request for the first 10 requests. After obtaining the anomaly scores, we separately compute the training and inference times. Training CoLLaTe on the MGAB dataset takes an average of 12.27 seconds per subset, while inference takes an average of 0.24 seconds per subset. We compare the running time of CoLLaTe and the baselines in detail in Table 6, where the time for LLM is calculated separately from the time for other steps. The memory size of the CoLLaTe (except the LLM part) is 43KB.
Regarding time complexity, the time complexity for transformer-based methods (LLM and TSADM in CoLLaTe) is $O(n^{2})$, where $n$ is the sequence length of prompt. The time complexity for the conditional network is $O(d^{2})$, where $d$ is the dimension of the feedforward network in the conditional network. Thus, the total time complexity of CoLLaTe is $O(n^{2}+d^{2})$.
A.15 Potential Limitation
Firstly, CoLLaTe is not suitable for application scenarios with limited computational capacity or unstable network accessibility, as we would be unable to deploy the LLM locally or call its API remotely. Secondly, CoLLaTe is not applicable in scenarios with high concurrency or stringent real-time requirements, as calling the GPT-4 API takes 1 second per request and triggers forced delays after several consecutive requests. Besides, if both LLM and TSADM in CoLLaTe fail to detect an anomaly, CoLLaTe will also fail to detect it.
A.16 Error Bar
We divide the time series by sliding window and each window consist of a data sample. We pick 50% of them as training set and 50% of them as test set. We divide the training set and test set by this method in ten different ways and train and test our model on them. Then, we compute the mean value and standard deviation (std) of CoLLaTe’ performance as the error bar, which is shown in Table. 7.
A.17 Assumption Validity
Assumptions 1 and 2 state that the anomaly score prediction errors of LLM and TSADM follow two unknown distributions, respectively. The only constraint on these distributions is that their expected values are not equal to zero. We consider this assumption to be practical, as no model can achieve perfect performance, and the expectation of their prediction errors should naturally bigger than zero. This is also proven by the experimental result shown in Tab. 2, where the F1 scores of CoLLaTe and GPT4 are not $1.0$. Assumption 3 assumes that $\mathcal{L}_{\beta}^{*}$ in Eq. 6 is $L$-Lipschitz continuous towards the trainable parameters in $\hat{S}_{i}$ and $\hat{S}_{j}$. We plot the variation of $|\mathcal{L}_{\beta}^{*}(\theta_{1})-\mathcal{L}_{\beta}^{*}(\theta_{2}|)$ as the distance between two trainable parameters $\theta_{1}$ and $\theta_{2}$ change in Fig. 5(d). We can assure $|\mathcal{L}_{\beta}^{*}(\theta_{1})-\mathcal{L}_{\beta}^{*}(\theta_{2}|)<L|\theta_{1}-\theta_{2}|$ when setting $L$ as $280$ and the Lipschitz continuous assumption is satisfied.
Table 6: The model efficiency.
Model Train (s/subset) Inference(s/subset) LLM (s/request)
CoLLaTe 12.27 0.24 1.27
GPT4 1.27
sigLLM 2.31
LLMAD 1.42
MSCRED 3.04 0.17
OmniAnomaly 2.32 0.14
TranAD 1.52 1.31
AnomalyTr 9.46 1.71
DCdetector 3.79 1.21
Table 7: Error bar of CoLLaTe.
Mustang MGAB Flight1 Flight2
Pre Rec F1 Pre Rec F1 Pre Rec F1 Pre Rec F1
Mean 0.950 1.000 0.971 0.975 0.996 0.985 0.790 0.978 0.866 0.897 0.891 0.883
Std. 0.112 0.000 0.064 0.054 0.012 0.030 0.138 0.044 0.079 0.094 0.048 0.062
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We accurately summarize the contribution and clarify the scope of CoLLaTe in abstract and introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitation is discussed in Appendix. A.15.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Assumptions are all listed in lines 176-180 and highlighted in bold. Proof is provided in Appendix. A.1, A.3, A.2, A.4.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The hyperparameter settings are provided in Table  5. Besides, the open-sourced datasets used in this paper and the code is provided in supplementary materials.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The open-sourced datasets used in this paper and the code is provided in supplementary materials.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All of these information can be found in our code, which will be released.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Error bar of CoLLaTe can be found in Appendix. A.16.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We list these information in Appendix. A.14.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We strictly conform the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: [N/A]
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: [N/A]
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We used public datasets and cited them as requested by the dataset authors.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We release them in supplementary material with readme.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: [N/A]
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: [N/A]
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: LLM is one of the components in our model.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.